# A Model Predictive Capture Point Control Framework for Robust Humanoid Balancing via Ankle, Hip, and Stepping Strategies

### v2

@@ Abstract
The robust balancing capability of humanoid robots has been considered one of the crucial requirements for their mobility in real environments. In particular, many studies have been devoted to the efficient implementation of human-inspired ankle, hip, and stepping strategies, to endow humanoids with human-level balancing capability. In this paper, a robust balance control framework for humanoids is proposed. Firstly, a Model Predictive Control (MPC) framework is proposed for Capture Point (CP) tracking control, enabling the integration of ankle, hip, and stepping strategies within a single framework. Additionally, a variable weighting method is introduced that adjusts the weighting parameters of the Centroidal Angular Momentum (CAM) damping control. Secondly, a hierarchical structure of the MPC and a stepping controller was proposed, allowing for the step time optimization. The robust balancing performance of the proposed method is validated through simulations and real robot experiments. Furthermore, a superior balancing performance is demonstrated compared to a state-of-the-art Quadratic Programming (QP)-based CP controller that employs the ankle, hip, and stepping strategies. The supplementary video is available at https://youtu.be/7Y4CykTpgrw
@@
Humanoid walking control, push recovery, capture point (CP) control, model predictive control (MPC), and three balance strategies (ankle, hip, and stepping strategies).
@@ I Introduction
Figure 1: Dynamic balancing of humanoid robot using ankle, hip, and stepping strategies in the presence of disturbances.
@@
In order to attain robust balancing performance in humanoid robots, researchers have been inspired by human balance strategies, and several studies are conducted to explore them [1, 2, 3]. The imitation and analysis of human balance strategies, including ankle, hip, and stepping strategies, have been explored using a simple model [4, 5, 6]. The lessons learned from these investigations have been applied to humanoid robot research, with each strategy being implemented in a distinct manner. Specifically, the ankle strategy has been implemented through the Zero Moment Point (ZMP) control and the hip strategy has been implemented using Centroidal Angular Momentum (CAM) control. Lastly, the stepping strategy has been realized through the stepping control, which adjusts footstep position or step time.
@@
In this paper, a robust balance control framework is proposed to overcome disturbances through the integration of the three balance strategies and step time optimization (as shown in Fig. 1). In contrast to [38, 39], our research does not restrict the horizon time of MPC based on the current step duration. The proposed framework expands upon our prior works [23, 31] that were developed for the same goal, i.e., CP tracking control. We combined the hip strategy in [23] and the ankle and stepping strategies in [31] into a single MPC framework. With novel ideas, we addressed several technical problems that arose during the integration process, thereby enhancing the coherence between control hierarchies in the proposed control framework. The primary contributions of this paper can be summarized as follows:
@@
    An MPC framework that integrates the three balance strategies for CP tracking control is proposed. The proposed MPC framework is an extension of the MPC framework proposed in our prior work [31]. Unlike [31], the proposed framework includes MPC optimization to handle not only ZMP control and stepping control but also CAM control. Furthermore, this framework enables more effective CAM control through MPC, unlike the heuristic CAM calculation method based on CMP decomposition proposed in our previous study [23].
@@
    A novel variable weighting method for CAM control is proposed. This method adjusts the weighting parameters of CAM damping control during the time horizon of the MPC to enhance the CP control performance and demonstrates better balancing performance compared to the conventional constant weighting method.
@@
    The proposed method is validated through extensive simulations and real robot experiments using our humanoid robot, TOCABI. Additionally, when compared to a state-of-the-art QP-based CP controller that incorporates the three balance strategies [39], the proposed method demonstrates superior balancing performance in the presence of disturbances.
This paper is organized as follows. In Section II, we briefly introduce the LIPFM and CP dynamics that underlie the walking control framework proposed in this paper. In Section III, an overview of the walking control framework is introduced. Section IV provides a detailed description of the MPC framework proposed in this paper. Section V describes the proposed MPC-based stepping controller. In Section VI, the results of simulations and real robot experiments conducted to validate the proposed algorithm are presented. In Section VII, discussions on the proposed framework and implementation details are provided. Finally, in Section VIII, the conclusion of this paper is presented. Subsequently, we shall refer to the proposed MPC framework as CP–MPC for the remainder of this paper.
@@ II-A Linear Inverted Pendulum Plus Flywheel Model
The dynamics of LIPFM in the y-direction can be derived in the same way as the x-direction, and the dynamics in each direction can be dealt with independently. This decoupling property of the system simplifies the analysis and control of the system and makes it easier to handle each direction separately. Therefore, in this study, the dynamics were derived only in the x-direction for conciseness of the paper, except for Section V, where the step time variable couples both x- and y-direction variables.
@@ III Overall Walking Control Structure
Figure 2: Overall control framework for robust humanoid walking; The three balance strategies are employed for CP tracking control based on the CP–MPC, and next footstep position and step time are determined by the stepping controller.
@@
The ZMP controller [14] computes desired contact wrenches to achieve the desired ZMP $\mathbf{z}^{des}=[z_{x}^{des}\,z_{y}^{des}]\in\mathbb{R}^{2}$, which is the first element of the ZMP control inputs $\mathbf{Z}$. Subsequently, the admittance controller compares the measured contact wrench $\mathbf{w}_{L,R}\in\mathbb{R}^{6}$ with the desired contact wrench to determine the changes of the position and orientation of the support foot with respect to the pelvis $\Delta\mathbf{e}_{L,R}\in\mathbb{R}^{6}$. The subscript $L$ and $R$ signify the left and right, respectively.
The stepping controller calculates the next footstep position $\mathbf{f}=[{f}_{x}\,{f}_{y}]\in\mathbb{R}^{2}$ and the step time $T$ to achieve the footstep adjustment $\Delta\mathbf{f}_{1}=[\Delta{f}_{x,1}\,\Delta{f}_{y,1}]\in\mathbb{R}^{2}$ planned by CP–MPC. Specifically, the controller primarily adjusts the step time $T$ based on CP end-of-step dynamics (5) to ensure that the $\mathbf{f}$ achieves $\Delta\mathbf{f}_{1}$ as closely as possible. Here, $\Delta\mathbf{f}_{1}$ refers to the first element of $\Delta\mathbf{F}$, representing the adjustment of the next footstep. After determining $\mathbf{f}$ and $T$, the foot trajectory generator creates reference foot trajectories $\mathbf{e}^{ref}_{L,R}\in\mathbb{R}^{6}$. Consequently, the lower body inverse kinematics is solved for the desired joint angles of legs $\mathbf{q}^{des}_{lb}$ required to achieve the reference CoM trajectories $\mathbf{c}^{ref}\in\mathbb{R}^{3}$, reference foot trajectories $\mathbf{e}^{ref}_{L,R}\in\mathbb{R}^{6}$, and displacement of the support foot pose $\Delta\mathbf{e}_{L,R}\in\mathbb{R}^{6}$. In addition, when a step change occurs, the footstep planner generates footsteps based on the target velocity command relative to the supporting foot.
The desired centroidal moment $\bm{\tau}^{des}=[\tau_{y}^{des}\,-\tau_{x}^{des}]\in\mathbb{R}^{2}$, which is the first element of the centroidal moment control inputs $\bm{\uptau}$ is controlled by the HQP-CAM controller proposed in our previous study [23]. The desired centroidal moment $\bm{\tau}^{des}$ is integrated to yield the desired CAM, and the HQP-CAM controller computes the desired upper body joint angles $\mathbf{q}^{des}_{ub}$ to achieve the desired CAM. In this study, only the CAM in the x- and y-directions is addressed based on the CP–CMP dynamics (4), while the CAM in the z-direction is not considered.
In the final stage, the joint PD controller is employed to calculate the joint torques required to track the upper and lower body motions. These computed joint torques $\bm{\Gamma}_{pd}$ are then commanded to the joints in conjunction with the gravity compensation torque $\bm{\Gamma}_{g}$.
@@ IV CP–MPC: Model Predictive Control Framework for Capture Point Tracking
In this section, we extend the framework of our previous study [31] to propose CP–MPC, an MPC framework that incorporates the three balance strategies: ankle, hip, and stepping strategies. The proposed CP–MPC utilizes not only the ZMP and footstep position but also the centroidal moment to enhance the CP control performance against strong external disturbances. Detailed explanations of the proposed MPC framework are provided in the following subsections.
@@ IV-A Prediction of Future Trajectory
where $A=e^{\omega T_{s}},\mathbf{B}=\left[1-e^{\omega T_{s}}\quad\frac{1-e^{\omega T_{s}}}{mg}\right]$. The CMP $p_{x}$ can be decomposed into ZMP $z_{x}$ and centroidal moment $\tau_{y}$ as shown in (4), and two variables are independently treated as control inputs of the CP–MPC.
By recursive application of (6), the predicted future trajectories of CP over the time horizon $N$ that emanate from the current CP $\xi_{x,k}$ can be expressed as below,
@@ IV-B Problem Setup for MPC Optimization
s. t. $\underline{\mathbf{Z}}_{x}\leq\mathbf{A}_{p}\begin{bmatrix}\mathbf{Z}_{x}\\ \Delta\mathbf{F}_{x}\end{bmatrix}\leq\overline{\mathbf{Z}}_{x}$ (11)
@@
${\xi^{ref}_{x,k+N}=\xi_{x,k+N}}$
with
$\mathbf{A}_{p}=\begin{bmatrix}\mathbf{I}_{N}&-\mathbf{S}\end{bmatrix},\quad\mathbf{S}=\begin{bmatrix}{s}_{1,1}&{s}_{1,2}&\cdots&{s}_{1,m}\\ \vdots&\vdots&\ddots&\vdots\\ {s}_{N,1}&{s}_{N,2}&\cdots&{s}_{N,m}\\ \end{bmatrix},$
${s_{k,m}=\left\{\begin{array}[]{ll}\text{{1; If time step $k$ is included in the $m$-th footstep}}\\ \text{{0; If time step $k$ is not included in the $m$-th footstep}}\end{array}\right.}$ (14)
@@
The cost function in (10) is composed of four cost terms in total. The first cost term functions as a reference CP trajectory tracking control and is weighted by a positive diagonal weighting matrix $\mathbf{w}_{\xi}\in\mathbb{R}^{N\times N}$. The second cost term plans the centroidal moment $\bm{\uptau}_{y}$ that drives the CAM $\mathbf{h}_{y}$ to zero when the disturbance is small. Here, $\mathbf{K}_{d}\in\mathbb{R}^{N\times N}$ is a damping matrix which is positive and diagonal and $\mathbf{h}_{y}\in\mathbb{R}^{N}$ is the CAM in y-direction which is an integration of centroidal moment $\bm{\uptau}_{y}$. This term is weighted by a positive diagonal matrix, $\mathbf{w}_{\tau}\in\mathbb{R}^{N\times N}$. The diagonal terms of $\mathbf{w}_{\tau}$ are adjusted based on the magnitude of the ZMP control inputs to improve the CP control performance. This adjustment algorithm is explained in Section IV-C. The third cost term regulates the additional footstep adjustments from the pre-planned footsteps. This term is weighted by a positive diagonal matrix $\mathbf{w}_{F}\in\mathbb{R}^{m\times m}$. The last cost term regulates the change of the control input to prevent rapid changes and generate a smooth control input. This term is weighted by positive diagonal matrix $\mathbf{w}_{p}\in\mathbb{R}^{2N\times 2N}$. The MPC variables consist of $\mathbf{P}_{x}$ and $\Delta\mathbf{F}_{x}$, where $\mathbf{P}_{x}\in\mathbb{R}^{2N}$ represents the sequence of ZMP $z_{x,k}$ and centroidal moment $\tau_{y,k}$, treated as control inputs, while $\Delta\mathbf{F}_{x}=[\Delta{f}_{x,1}\Delta{f}_{x,2}\cdots\Delta{f}_{x,m}]^{\textrm{T}}\in\mathbb{R}^{m}$ represents a vector consisting of additional footstep adjustments in the x-direction.
@@
The constraints in (11) consist of three inequality conditions and one equality condition. The first constraint refers to the ZMP constraint, which confines the ZMP control input $\mathbf{Z}_{x}\in\mathbb{R}^{N}$ within the support polygon. The vectors $\overline{\mathbf{Z}}_{x}\in\mathbb{R}^{N}$ and $\underline{\mathbf{Z}}_{x}\in\mathbb{R}^{N}$ indicate the upper and lower limits of ZMP inputs, respectively, and these limits are determined by reflecting the size of both feet approximated as rectangles based on the reference ZMP trajectory $\mathbf{Z}^{ref}$. In the matrix $\mathbf{A}_{p}$, the matrix $\mathbf{I}_{N}\in\mathbb{R}^{N\times N}$ refers to an $N$ size identity matrix and the matrix $\mathbf{S}\in\mathbb{R}^{N\times m}$ represents a selection matrix whose element ${s}_{k,m}$ is 1 or 0. The element ${s}_{k,m}$ is set to 1 to activate $\Delta{f}_{x,m}$ when time step $k$ is included in the $m$-th footstep after the current Single Support Phase (SSP), and it is set to 0 when time step $k$ is not included in the $m$-th footstep, including the current SSP. Using $\Delta\mathbf{F}_{x}$ to relax the ZMP constraint in situations where ZMP control inputs are limited improves the performance of CP control by increasing flexibility in generating ZMP control inputs. The second constraint refers to the limit on the amount of centroidal moment that can be produced by the robot’s motion. The vectors $\overline{\bm{\uptau}}_{y}\in\mathbb{R}^{N}$ and $\underline{\bm{\uptau}}_{y}\in\mathbb{R}^{N}$ indicate the upper and lower bounds of the centroidal moment, respectively. The third constraint denotes the kinematic constraint of the additional footsteps to prevent singularities and self-collisions between legs. The vectors $\Delta\overline{\mathbf{F}}_{x}\in\mathbb{R}^{m}$ and $\Delta\underline{\mathbf{F}}_{x}\in\mathbb{R}^{m}$ refer to the upper and lower bounds of additional footsteps adjustment. The last constraint signifies the terminal constraint imposed to ensure the convergence of CP tracking.
@@
TABLE I: Parameters of CP–MPC and stepping controller used in the simulations and real robot experiments.

### v3

@@ Abstract
The robust balancing capability of humanoids is essential for mobility in real environments. Many studies focus on implementing human-inspired ankle, hip, and stepping strategies to achieve human-level balance. In this paper, a robust balance control framework for humanoids is proposed. Firstly, a Model Predictive Control (MPC) framework is proposed for Capture Point (CP) tracking control, enabling the integration of ankle, hip, and stepping strategies within a single framework. Additionally, a variable weighting method is introduced that adjusts the weighting parameters of the Centroidal Angular Momentum damping control. Secondly, a hierarchical structure of the MPC and a stepping controller was proposed, allowing for the step time optimization. The robust balancing performance of the proposed method is validated through simulations and real robot experiments. Furthermore, a superior balancing performance is demonstrated compared to a state-of-the-art Quadratic Programming-based CP controller that employs the ankle, hip, and stepping strategies.
@@
Humanoid walking control, capture point (CP) control, and model predictive control (MPC).
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
In order to obtain robust balancing performance in humanoid robots, researchers have been inspired by human balance strategies, and several studies are conducted to explore them [1, 2, 3]. The imitation and analysis of human balance strategies, including ankle, hip, and stepping strategies, have been explored using a simple model [4, 5, 6]. The lessons learned from these investigations have been applied to humanoid robot research, with each strategy being implemented in a distinct manner. Specifically, the ankle strategy has been implemented through the Zero Moment Point (ZMP) control and the hip strategy has been implemented using Centroidal Angular Momentum (CAM) control. Lastly, the stepping strategy has been realized through the stepping control, which adjusts footstep position or step time.
@@
In this paper, a robust balance control framework is proposed to overcome disturbances through the integration of the three balance strategies and step time optimization. In contrast to [38, 39], our research does not restrict the horizon time of MPC based on the current step duration. The proposed framework expands upon our prior works [23, 31] that were developed for the same goal, i.e., CP tracking control. We combined the hip strategy in [23] and the ankle and stepping strategies in [31] into a single MPC framework. With novel ideas, we addressed several technical problems that arose during the integration process, thereby enhancing the coherence between control hierarchies in the proposed control framework. The primary contributions of this paper can be summarized as follows:
@@
    An MPC framework that integrates three balance strategies for CP tracking control is proposed as an extension of our prior work [31]. Unlike [31], this framework optimizes not only ZMP and stepping control but also CAM control. Additionally, it enhances CAM control through MPC, rather than using the heuristic CAM calculation method proposed in our previous study [23].
@@
    A novel variable weighting method for CAM control is proposed. This method adjusts the weighting parameters of CAM damping control during the time horizon of the MPC to enhance the CP control performance.
@@
    A novel HQP-based whole-body IK controller is proposed, focusing on nominal pose tracking and CAM control, while ensuring a smooth transition into initial pose recovery.
  5. 5.
    The proposed method is validated through extensive simulations and real robot experiments, showing superior balancing performance compared to a state-of-the-art QP-based CP controller [39] that incorporates the three balance strategies.
This paper is organized as follows. Section II introduces the LIPFM and CP dynamics, followed by an overview of the walking control framework in Section III. The proposed MPC framework is detailed in Section IV, and the MPC-based stepping controller is described in Section V. Section VI presents the HQP-based WBIK. Simulation and experimental results validating the proposed algorithm are provided in Section VII, while Section VIII discusses the framework and implementation details. Finally, the conclusion is presented in Section IX. Henceforth, the proposed MPC framework is referred to as CP–MPC.
@@ II-A Linear Inverted Pendulum Plus Flywheel Model
The dynamics of LIPFM in the y-direction can be derived in the same way as the x-direction, and the dynamics in each direction can be dealt with independently. Therefore, in this study, the dynamics were derived only in the x-direction for conciseness of the paper, except for Section V, where the step time variable couples both x- and y-direction variables.
@@ III Overall Walking Control Structure
Figure 1: Overall walking control framework: Three balance strategies for CP tracking control via CP–MPC, footstep position and step time determination by the stepping controller, and implementation through HQP–WBIK.
@@
The ZMP controller [14] computes desired contact wrenches to achieve the desired ZMP $\mathbf{z}^{des}=[z_{x}^{des}\,z_{y}^{des}]\in\mathbb{R}^{2}$, which is the first element of the ZMP control inputs $\mathbf{Z}$. Subsequently, the admittance controller compares the measured contact wrench $\mathbf{w}_{L,R}\in\mathbb{R}^{6}$ with the desired contact wrench to determine the changes in the position and orientation of the support foot with respect to the pelvis $\Delta\mathbf{e}_{L,R}\in\mathbb{R}^{6}$, where both wrenches reflect only the z-direction position and roll/pitch orientation components. The subscript $L$ and $R$ signify the left and right, respectively.
The stepping controller calculates the next footstep position $\mathbf{f}=[{f}_{x}\,{f}_{y}]\in\mathbb{R}^{2}$ and the step time $T$ to achieve the footstep adjustment $\Delta\mathbf{f}_{1}=[\Delta{f}_{x,1}\,\Delta{f}_{y,1}]\in\mathbb{R}^{2}$ planned by CP–MPC. Specifically, the controller primarily adjusts the step time $T$ based on CP end-of-step dynamics (5) to ensure that the $\mathbf{f}$ achieves $\Delta\mathbf{f}_{1}$ as closely as possible. Here, $\Delta\mathbf{f}_{1}$ refers to the first element of $\Delta\mathbf{F}$, representing the adjustment of the next footstep. After determining $\mathbf{f}$ and $T$, the foot trajectory generator creates reference foot trajectories $\mathbf{e}^{ref}_{L,R}\in\mathbb{R}^{6}$. In addition, when a step change occurs, the footstep planner generates footsteps based on the target velocity command relative to the supporting foot.
The desired centroidal moment $\bm{\tau}^{des}=[\tau_{y}^{des},-\tau_{x}^{des}]\in\mathbb{R}^{2}$, which is the first element of the centroidal moment control input $\bm{\uptau}$, is integrated to obtain the desired CAM. In this study, only the CAM in the x- and y-directions is computed from the CP–MPC based on CP–CMP dynamics (4), while the CAM in the z-direction is controlled to remain zero to compensate for momentum induced by the swing foot motion and to prevent yaw slip at the support foot.
In the final stage, the HQP-based whole-body IK receives feedback from the FK results, represented as $\mathbf{X}_{FK}=[\mathbf{c}~{}\mathbf{e}~{}\mathbf{x}_{pelv}~{}\mathbf{x}_{chest}~{}\mathbf{x}_{hand}]^{T}$, and determines the desired joint angles $\mathbf{q}^{des}$ and joint velocities $\mathbf{\dot{q}}^{des}$ required to achieve the reference CoM trajectory $\mathbf{c}^{ref}\in\mathbb{R}^{3}$, reference foot trajectories $\mathbf{e}^{ref}_{L,R}\in\mathbb{R}^{6}$, support foot pose displacement $\Delta\mathbf{e}_{L,R}\in\mathbb{R}^{6}$, and desired CAM $\mathbf{h}^{des}\in\mathbb{R}^{3}$. Then, a joint PD controller computes the torques $\bm{\Gamma}_{pd}$ to track the desired joint trajectory, applied alongside the gravity compensation torque $\bm{\Gamma}_{g}$ computed using the contact-consistent control framework [45].
@@ IV CP–MPC: Model Predictive Control Framework for Capture Point Tracking
In this section, we extend our previous framework [31] to propose CP–MPC, an MPC framework integrating ankle, hip, and stepping strategies for balance control. CP–MPC leverages the ZMP, footstep position, and centroidal moment to enhance CP control under strong external disturbances. The following subsections provide a detailed description of the proposed framework.
@@ IV-A Prediction of Future Trajectory
where $A=e^{\omega T_{s}},\mathbf{B}=\left[1-e^{\omega T_{s}}\quad\frac{1-e^{\omega T_{s}}}{mg}\right]$. The CMP $p_{x}$ can be decomposed into ZMP $z_{x}$ and centroidal moment $\tau_{y}$ as shown in (4), and two variables are independently treated as control inputs of the CP–MPC. By recursive application of (6), the predicted future trajectories of CP over the time horizon $N$ that emanate from the current CP $\xi_{x,k}$ can be expressed as below,
@@ IV-B Problem Setup for MPC Optimization
s. t. $\underline{\mathbf{Z}}_{x}\leq\begin{bmatrix}\mathbf{I}_{N}&-\mathbf{S}\end{bmatrix}\begin{bmatrix}\mathbf{Z}_{x}\\ \Delta\mathbf{F}_{x}\end{bmatrix}\leq\overline{\mathbf{Z}}_{x}$ (11)
@@
${\xi^{ref}_{x,k+N}=\xi_{x,k+N}.}$
@@
The cost function in (10) is composed of four cost terms in total. The first cost term functions as a reference CP trajectory tracking control and is weighted by a positive diagonal weighting matrix $\mathbf{w}_{\xi}\in\mathbb{R}^{N\times N}$. The second cost term plans the centroidal moment $\bm{\uptau}_{y}$ that drives the CAM $\mathbf{h}_{y}$ to zero when the disturbance is small. Here, $\mathbf{K}_{d}\in\mathbb{R}^{N\times N}$ is a damping matrix which is positive and diagonal and $\mathbf{h}_{y}\in\mathbb{R}^{N}$ is the CAM in y-direction which is an integration of centroidal moment $\bm{\uptau}_{y}$. This term is weighted by a positive diagonal matrix, $\mathbf{w}_{\tau}\in\mathbb{R}^{N\times N}$. The diagonal terms of $\mathbf{w}_{\tau}$ are adjusted based on the magnitude of the ZMP control inputs to improve the CP control performance. This adjustment algorithm is explained in Section IV-C. The third cost term regulates the additional footstep adjustments from the pre-planned footsteps. This term is weighted by a positive diagonal matrix $\mathbf{w}_{F}\in\mathbb{R}^{m\times m}$. The last cost term regulates the change of the control input to prevent rapid changes and generate a smooth control input. This term is weighted by positive diagonal matrix $\mathbf{w}_{p}\in\mathbb{R}^{2N\times 2N}$. The MPC variables consist of $\mathbf{P}_{x}$ and $\Delta\mathbf{F}_{x}$, where $\mathbf{P}_{x}\in\mathbb{R}^{2N}$ represents the sequence of ZMP $z_{x,k}$ and centroidal moment $\tau_{y,k}$, treated as control inputs, while $\Delta\mathbf{F}_{x}=[\Delta{f}_{x,1}\Delta{f}_{x,2}\cdots\Delta{f}_{x,m}]^{\textrm{T}}\in\mathbb{R}^{m}$ represents a vector consisting of additional footstep adjustments in the x-direction. Here, although $\Delta\mathbf{F}_{x}$ is not defined as a variable in CP–CMP dynamics (6), it is utilized as an optimization variable to relax the ZMP constraint.
@@
The constraints in (11) consist of three inequality conditions and one equality condition. The first constraint refers to the ZMP constraint, which confines the ZMP control input $\mathbf{Z}_{x}\in\mathbb{R}^{N}$ within the support polygon. The vectors $\overline{\mathbf{Z}}_{x}\in\mathbb{R}^{N}$ and $\underline{\mathbf{Z}}_{x}\in\mathbb{R}^{N}$ indicate the upper and lower limits of ZMP inputs, respectively. These limits are approximated by modeling the foot as a rectangle, defined by the foot’s width and length relative to the reference ZMP $\mathbf{Z}^{ref}$, while taking yaw rotation into account. To prevent the desired ZMP from being generated outside the support polygon due to approximation errors, we approximate the rectangular foot size such that the approximated support area falls within a range smaller than the actual foot shape-based support area. The matrix $\mathbf{I}_{N}\in\mathbb{R}^{N\times N}$ refers to an $N$ size identity matrix and the matrix $\mathbf{S}\in\mathbb{R}^{N\times m}$ represents a selection matrix whose element ${s}_{k,m}$ is 1 or 0. The element ${s}_{k,m}$ is set to 1 to activate $\Delta{f}_{x,m}$ when time step $k$ is included in the $m$-th footstep after the current Single Support Phase (SSP), and it is set to 0 when time step $k$ is not included in the $m$-th footstep, including the current SSP. Using $\Delta\mathbf{F}_{x}$ to relax the ZMP constraint in situations where ZMP control inputs are limited improves the performance of CP control by increasing flexibility in generating ZMP control inputs. The second constraint refers to the limit on the amount of centroidal moment that can be produced by the robot’s motion. The vectors $\overline{\bm{\uptau}}_{y}\in\mathbb{R}^{N}$ and $\underline{\bm{\uptau}}_{y}\in\mathbb{R}^{N}$ indicate the upper and lower bounds of the centroidal moment, respectively. These constraints were experimentally determined by considering joint position and velocity limits. The third constraint denotes the kinematic constraint of the additional footsteps to prevent singularities and self-collisions between legs. The vectors $\Delta\overline{\mathbf{F}}_{x}\in\mathbb{R}^{m}$ and $\Delta\underline{\mathbf{F}}_{x}\in\mathbb{R}^{m}$ refer to the upper and lower bounds of additional footsteps adjustment. The last constraint signifies the terminal constraint imposed to ensure the convergence of CP tracking.
@@
TABLE I: Parameters used in the controllers for simulations and real robot experiments.
@@
HQP–WBIK
Parameters Value
$\mathbf{w}_{J}$
$120.0$ (Foot pose), $10.0$ (CoM), $0.1$ (Hand pose),
$1.0$ (Chest orientation), $5.0$ (Pelvis orientation)
$\mathbf{w}_{A}$ $1.0$
$\mathbf{w}_{\dot{q}}$ $0.0001$
$\mathbf{K}_{p}$
$30.0$ (Foot pose), $30.0$ (CoM), $10.0$ (Hand pose),
$10.0$ (Chest orientation), $10.0$ (Pelvis orientation)
$\mathbf{K}_{q}$ $50.0$